3i)16 in polar form: z r(cos 0isin 0) where (1 Write the complex number z and e= The angle should satisfy 0 0 < 2«.
Answer #1

7=61-13 )! {t-1+53:) 27 - {1, 2(-1) (131) +(81) = (1 - 27372-3) 6-2-253138 = (-238 (1+131) 8 : 256 {(1473 ;)) -256 (142631- 3if you faced any problem to understand it contact me on comment section.